OpenStudy (destinyyyy):

Subtract and simplify answer. 3x/x^2+4x+3 - x-6/x+1

OpenStudy (destinyyyy):

I got the answer -x^2-9x-18/(x+3)(x+1) .. Its wrong. -6x is the correct part not -9x... How is it -6x?

OpenStudy (anonymous):

this? \[\frac{ 3x }{ x^2+4x+3 }-\frac{ x-6 }{ x+1 }\]

OpenStudy (destinyyyy):

Yes.

OpenStudy (anonymous):

factor on the right and get a common denominator \[\frac{ 3x }{ (x+3)(x+1) }-\frac{ (x-6)(x+3) }{ (x+3)(x+1) }\] combine into one fraction \[\frac{ 3x-((x-6)(x+3)) }{ (x+3)(x+1) }\] foil \[\frac{ 3x-(x^2-6x+3x-18) }{ (x+3)(x+1) }\] distribute the negative \[\frac{ 3x-x^2+6x-3x+18 }{ (x+3)(x+1) }\] combine like terms \[\frac{ -x^2+6x+18 }{ (x+3)(x+1) }\]

OpenStudy (anonymous):

or if you factor out the negative in front of x² \[-\frac{ x^2-6x-18 }{ (x+3)(x+1) }\]

OpenStudy (destinyyyy):

Okay I see how you got 6.. But why did you put -6 before 3? I have --- (x-6)(x+3)== (x^2+3x-6x-18)

OpenStudy (anonymous):

addition is commutative, meaning the order doesn't matter

OpenStudy (anonymous):

1 + 2 is the same as 2 + 1

OpenStudy (destinyyyy):

Okay..

OpenStudy (destinyyyy):

What your saying isnt right.. Your answer-> 3+6-3=6 Mine-->3+3-6=0 By changing where the numbers go gives a completely different number.

OpenStudy (destinyyyy):

Never mind Ill wait for a response from my teacher. Thanks for I guess helping.

OpenStudy (anonymous):

I wouldn't know what yours is, or how you got it because you haven't shown me any work.

OpenStudy (destinyyyy):

All you had to do was ask.. x^2+4x+3= (x+3)(x+1) 3x/ (x+3)(x+1) - x-6/x+1 * (x+3)/(x+3) 3x-(x-6)(x+3)/ (x+3)(x+1) 3x-(x^2+3x-6x-18/ (x+3)(x+1) --3x-6x=-3-- -x^2-9x-18/(x+3)(x+1)

OpenStudy (anonymous):

no offense, but you're asking the question. I really shouldn't have to ask to see your work. This is your 4th line: 3x-(x^2+3x-6x-18)/ (x+3)(x+1) You have to distribute the - sign. It's kind of like a -1 in front of the parentheses. 3x -1(x² + 3x - 6x - 18) 3x - x² - 3x + 6x + 18 Just the x terms is 3x - 3x + 6x so the whole numerator is -x² + 6x + 18

OpenStudy (anonymous):

it just looks like you didn't distribute